OpenStudy (anonymous):

Evaluate 5! + 2!. a. 60 b. 122 c. 5,040

OpenStudy (whpalmer4):

5! = 5*4*3*2*1 2! = 2*1 n! = n*(n-1)*...*2*1

OpenStudy (anonymous):

so, the exclamation point means multiply the number by all the numbers below it?

OpenStudy (whpalmer4):

yeah, it's called the factorial operator.

OpenStudy (anonymous):

ok thanks